# Risk premium and rough volatility

### v1

@@ Abstract.
60F17, 60F05, 60G15, 60G22, 91G20, 91G60, 91B25
OB and AJ are supported by the EPSRC grant EP/T032146/1. For the purpose of open access, the author(s) has applied a Creative Commons Attribution (CC BY) licence (where permitted by UKRI, ‘Open Government Licence’ or ‘Creative Commons Attribution No-derivatives (CC BY-ND) licence’ may be stated instead) to any Author Accepted Manuscript version arising’
@@ Introduction
Rough volatility is a recent paradigm proposed by Gatheral, Jaisson and Rosenbaum [10], which has attracted the attention of many academics and practitioners thanks to its numerous attractive properties. Despite some debate about whether volatility should be rough [12, 13, 17, 16, 23], this class of models provides a general framework to analyse both time series of the instantaneous volatility (under the historical measure $\mathbb{P}$) and prices of financial derivatives (under the pricing measure $\mathbb{Q}$). Starting from a rough version of the Bergomi model [5] under $\mathbb{P}$, Bayer, Friz and Gatheral [3] showed that a deterministic market price of risk preserved its structure under $\mathbb{Q}$ (somehow akin to the Heston model [14] specification).
However, the Financial Economics literature has long shown that this market price of risk, monitoring the transition from $\mathbb{P}$ to $\mathbb{Q}$ via Girsanov’s transform, is not constant nor deterministic but instead stochastic. Its estimation has been the source of long academic discussions, outside the scope of the present paper though, and we refer the interested reader to [1, 6, 7, 8, 19, 21] for some useful pointers. This of course has serious practicals implications for risk management, and [11, 24] are fascinating sources of information. We focus here on this particular bridge between $\mathbb{P}$ and $\mathbb{Q}$ and show, not surprisingly, that the required stochasticity of the market price of risk unfortunately breaks the structure of the rough Bergomi model under $\mathbb{Q}$. However, we link the Hölder regularity of the volatility process (lower in this class of rough models) with that of the change of measure, and design several specifications making the model tractable under $\mathbb{Q}$. While the rough Bergomi model tracks well the behaviour of the historical volatility, it is less powerful for option prices, especially when considering VIX smiles (which are more or less flat under this model). Our new setup allows for more flexibility there, while preserving the $\mathbb{P}$-tractability of the model.
@@ 1. Rough volatility models and change of measure
Rough volatility models are a natural extension of classical stochastic volatility models. Starting from such a model under the historical measure $\mathbb{P}$, we characterise below its dynamics under equivalent martingale measures $\mathbb{Q}\sim\mathbb{P}$, which then, by the fundamental theorem of asset pricing, allows for arbitrage-free option pricing. Following for example [4, 10] we consider a rather general class of (rough) stochastic volatility model under $\mathbb{P}$, where the stock price process admits the following dynamics:
@@
starting from $S_{0}>0$, over a fixed time interval $\mathbb{T}:=[0,T]$, for $T>0$. Here $\boldsymbol{W}^{\mathbb{P}}=(W^{\mathbb{P}},W^{\mathbb{P},\perp})$ is a two-dimensional standard Brownian motion defined on a given filtered probability space $(\Omega,\mathcal{F},\mathbb{P})$, with $\mathcal{F}=\mathcal{F}^{W^{\mathbb{P}}}\vee\mathcal{F}^{W^{\mathbb{P},\perp}}$, and $Z^{\mathbb{P}}:=\rho W^{\mathbb{P}}+\overline{\rho}W^{\mathbb{P},\perp}$, with $\rho\in[-1,1]$ and $\overline{\rho}:=\sqrt{1-\rho^{2}}$. We further introduce the set $\mathbb{F}_{b}$ of $\mathbb{P}$-bounded and $(\mathcal{F}_{t})_{t\in\mathbb{T}}$-progressively measurable process and recall the Doléans-Dade stochastic exponential of a square integrable process $X$:
$\mathcal{E}(X)_{t}:=\exp{\left(X_{t}-\frac{1}{2}\langle X\rangle_{t}\right)},\qquad t\in\mathbb{T}.$
@@
    For each $t\in\mathbb{T}$, $\mathbb{E}\left[v_{t}^{-1}\right]$ is finite;
@@
    Given an interest rate process $r:=(r_{t})_{t\in\mathbb{T}}\in\mathbb{F}_{b}$, there exist a sequence (^$(\widehat{\mathbb{P}}_{n})_{n\in\mathbb{N}}$ and a process $\gamma:=(\gamma_{t})_{t\in\mathbb{T}}$ in $\mathbb{F}_{b}$ and bounded ^$\widehat{\mathbb{P}}_{n}$-almost surely, for each $n\in\mathbb{N}$, such that
    $\sup_{t\in\mathbb{T}}\left\{\rho\int_{0}^{t}k(t,u)\chi_{u}\mathrm{d}u\right\}\leq 0,\qquad\widehat{\mathbb{P}}_{n}\text{-almost surely},\text{ for all }n\in\mathbb{N},$
    where we introduce the Sharpe ratio $\chi_{u}:=\frac{r_{u}-\mu_{u}}{\sqrt{v_{u}}}$ and
    (1.2) $\frac{\mathrm{d}\widehat{\mathbb{P}}_{n}}{\mathrm{d}\mathbb{P}}\bigg{|}_{\mathcal{F}_{t}}:=\mathcal{E}\left(\int_{0}^{\cdot}\chi_{u}\mathrm{d}W^{\mathbb{P}}_{u}+\int_{0}^{\cdot}\gamma_{u}\mathrm{d}W^{\mathbb{P},{\perp}}_{u}\right)_{t\wedge\tau_{n}}\quad\text{and}\quad\tau_{n}:=\inf\{t\geq 0,\;Y_{t}=n\};$
  5. (v)
    The correlation is negative: $\rho\leq 0$;
  6. (vi)
    For each $t\in\mathbb{T}$, the kernel $k(t,\cdot)$ is null on $\mathbb{T}\setminus[0,t]$ and $\int_{0}^{\cdot}k(\cdot,s)\mathrm{d}Z^{\mathbb{P}}_{s}$ is a well-defined Gaussian process.
@@
[…]
The following examples are common choices of such kernels:
Example 1.3.
The Gamma kernel, common in the $\mathcal{BSS}$ literature pioneered by Barndorff-Nielsen and Schmiegel [2], is given by
(1.3) $k(t,s)=(t-s)^{H-\frac{1}{2}}\mathrm{e}^{-\beta(t-s)}1{1}_{\{t\geq s\}},\qquad\text{with }H\in(0,1),\quad\beta\geq 0.$
(1.4) $\mathcal{D}^{\gamma}_{t}:=\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}\bigg{|}_{\mathcal{F}_{t}^{\mathbb{P}}}=\mathcal{E}\left(\int_{0}^{\cdot}\chi_{u}\mathrm{d}W^{\mathbb{P}}_{u}+\int_{0}^{\cdot}\gamma_{u}\mathrm{d}W^{\mathbb{P},{\perp}}_{u}\right)_{t},$
so that, from (1.2) above, $\mathcal{D}^{\gamma}_{t\wedge\tau_{n}}=\frac{\mathrm{d}\widehat{\mathbb{P}}_{n}}{\mathrm{d}\mathbb{P}}\bigg{|}_{\mathcal{F}_{t}}$.
Proposition 1.4.
For any $\gamma$ satisfying Assumption 1.1(iv), the process $\int_{0}^{\cdot}\chi_{u}\mathrm{d}W^{\mathbb{P}}_{u}+\int_{0}^{\cdot}\gamma_{u}\mathrm{d}W^{\mathbb{P},{\perp}}_{u}$ is a $\mathbb{P}$-local martingale on $\mathbb{T}$.
Proof.
Recall that the sum of two local martingale (with respect to the same filtration) is a local martingale whose sequence of stopping times is given by the minimum of the sequences of stopping times for the terms in the sum. The boundedness of $\gamma$ implies that $\int_{0}^{\cdot}\gamma_{s}\mathrm{d}W^{\mathbb{P},{\perp}}_{s}$ is a true $\mathbb{P}$-martingale. Regarding $\int_{0}^{\cdot}\chi_{u}\mathrm{d}W^{\mathbb{P}}_{u}$, the integrand is $\mathcal{F}_{s}$-measurable and locally bounded with respect to the sequence of stopping times $(\tau_{n})_{n\in\mathbb{N}}$ in (1.2) thanks to Assumption 1.1(i)-(iii)-(iv). Exploiting Assumption 1.1(ii)-(iii)-(iv), then $\mathbb{E}^{\mathbb{P}}[\chi_{t}^{2}]$ is finite for all $t\in\mathbb{T}$. Thus, the first term in the sum is a local martingale as well, and so is the whole sum. ∎
Proposition 1.4 justifies the use of a Doléans-Dade stochastic exponential in the definition of $\mathcal{D}^{\gamma}$. Furthermore, we obtain the following result.
@@
Under Assumption 1.1, for any $\gamma$ as in Assumption 1.1(iv), then
@@
    with $S_{0},v_{0}>0$, and $\lambda$ is the market price of volatility risk defined by
    (1.6) $\lambda_{t}:=\rho\chi_{t}+\overline{\rho}\;\gamma_{t},$
    and where $W^{\mathbb{Q}}$ and $Z^{\mathbb{Q}}$ are $\mathbb{Q}$-Brownian motions defined as
[…]
@@
    the discounted stock price $\widetilde{S}_{t}:=\frac{S_{t}}{B_{t}}$ with $\mathrm{d}B_{t}=r_{t}B_{t}\mathrm{d}t$, $B_{0}=1$, is a true $\mathbb{Q}$-martingale.
@@
To satisfy the no-arbitrage conditions, the change of measure for $W^{\mathbb{P}}$ is constrained by the martingale restriction on the discounted spot dynamics, while the Brownian motion $Z^{\mathbb{P}}$ gives freedom to the model and makes the market incomplete by the free choice of the process $\gamma$. Consequently, the change of measure from $\mathbb{P}$ to $\mathbb{Q}$ and the corresponding Radon-Nikodym derivative directly follow from Girsanov’s Theorem via (1.4), provided that $\mathcal{D}^{\gamma}_{t}\in L^{1}$ and $\mathcal{D}^{\gamma}$ is a true martingale. Thus, once we have shown (I), then (II) automatically follows. By Proposition 1.4, $\mathcal{D}^{\gamma}_{t}\in L^{1}$ and, being a non-negative local martingale, it is a supermartingale, and a true martingale on $\mathbb{T}$ if and only if $\mathbb{E}\left[\mathcal{D}^{\gamma}_{T}\right]=1$. To prove this, we closely follow [9, Proof of Theorem 1.1] with some modifications, and recall the stopping time $\tau_{n}:=\inf\{t\geq 0,\;Y_{t}=n\}$ from (1.2). For any $s\in\mathbb{T}$, the random function $f(x):=\frac{r_{s}-\mu_{s}}{\sqrt{\psi(s,x)}}$ is $\mathbb{P}$-bounded on $(-\infty,a]$ for any $a>0$ since $r$ and $\mu$ are $\mathbb{P}$-bounded, and $\psi(s,\cdot)$ bounded away from zero on intervals of the form $(-\infty,a]$, by Assumption 1.1(iii)-(iv)-(i), respectively. Then, again by Proposition 1.4,
(1.8) $1=\mathbb{E}\left[\mathcal{D}^{\gamma}_{T\wedge\tau_{n}}\right]=\mathbb{E}\left[\mathcal{D}^{\gamma}_{T}1{1}_{\{T<\tau_{n}\}}\right]+\mathbb{E}\left[\mathcal{D}^{\gamma}_{\tau_{n}}1{1}_{\{\tau_{n}\leq T\}}\right].$
The first term in (1.8) converges to $\mathbb{E}\left[\mathcal{D}^{\gamma}_{T}\right]$ as $n$ tends to infinity, yielding
$1-\mathbb{E}\left[\mathcal{D}^{\gamma}_{T}\right]=\lim_{n\uparrow\infty}\mathbb{E}\left[\mathcal{D}^{\gamma}_{\tau_{n}}1{1}_{\{\tau_{n}\leq T\}}\right].$
Girsanov’s theorem implies $\mathbb{E}\left[\mathcal{D}^{\gamma}_{\tau_{n}}1{1}_{\{\tau_{n}\leq T\}}\right]=\widehat{\mathbb{P}}_{n}(\tau_{n}\leq T)$, where ^$\widehat{\mathbb{P}}_{n}$ is defined such that
$\widehat{W}^{n}_{t}=\displaystyle W^{\mathbb{P}}_{t}-\int_{0}^{t\wedge\tau_{n}}\chi_{u}\mathrm{d}u$
is a ^$\widehat{\mathbb{P}}_{n}$-Brownian motion. Then, under ^$\widehat{\mathbb{P}}_{n}$, the process $Y$ becomes
$Y_{t}=\widehat{Y}_{t}^{n}+\int_{0}^{t\wedge\tau_{n}}k(t,s)\Big{(}\rho\chi_{u}+\overline{\rho}\;\gamma_{u}\Big{)}\mathrm{d}u,$
where $\widehat{Y}^{n}_{t}:=\int_{0}^{t}k(t,s)\mathrm{d}\widehat{Z}^{n}_{s}$ and
$\widehat{Z}^{n}_{t}=Z^{\mathbb{P}}_{t}-\int_{0}^{t\wedge\tau_{n}}\Big{(}\rho\chi_{u}+\overline{\rho}\;\gamma_{u}\Big{)}\mathrm{d}u,$
for $t\geq 0$, where $\widehat{Z}^{n}$ is a ^$\widehat{\mathbb{P}}_{n}$-Brownian motion. Furthermore, by Assumption 1.1(iv), there exists $M_{\gamma}>0$ such that $|\gamma_{t}|<M_{\gamma}$$\widehat{\mathbb{P}}_{n}$-almost surely, and
^$\widehat{\mathbb{P}}_{n}\left(\sup_{t\in\mathbb{T}}Y_{t}\geq n\right)$ =^$=\widehat{\mathbb{P}}_{n}\left(\sup_{t\in\mathbb{T}}\left\{\widehat{Y}_{t}^{n}+\int_{0}^{t\wedge\tau_{n}}k(t,u)\Big{(}\rho\chi_{u}+\overline{\rho}\;\gamma_{u}\Big{)}\mathrm{d}s\right\}\geq n\right)$
^$\leq\widehat{\mathbb{P}}_{n}\left(\sup_{t\in\mathbb{T}}\widehat{Y}_{t}^{n}+\sup_{t\in\mathbb{T}}\left\{\rho\int_{0}^{t\wedge\tau_{n}}k(t,u)\chi_{u}\mathrm{d}u\right\}+\sup_{t\in\mathbb{T}}\left\{\overline{\rho}\int_{0}^{t\wedge\tau_{n}}k(t,u)\gamma_{u}\mathrm{d}u\right\}\geq n\right)$
^$\leq\widehat{\mathbb{P}}_{n}\left(\sup_{t\in\mathbb{T}}\widehat{Y}_{t}^{n}+\sup_{t\in\mathbb{T}}\left\{\rho\int_{0}^{t\wedge\tau_{n}}k(t,u)\chi_{u}\mathrm{d}u\right\}\geq n-|\overline{\rho}|K_{\mathbb{T}}M_{\gamma}\right),$
where $K_{\mathbb{T}}:=\sup_{t\in\mathbb{T}}\int_{0}^{t}k(t,s)\mathrm{d}s$. Then, an application of Assumption 1.1(iv), we obtain
(1.9) ^$\widehat{\mathbb{P}}_{n}\left(\sup_{t\in\mathbb{T}}Y_{t}\geq n\right)\leq\widehat{\mathbb{P}}_{n}\left(\sup_{t\in\mathbb{T}}\widehat{Y}_{t}^{n}\geq n-|\overline{\rho}|K_{\mathbb{T}}M_{\gamma}\right).$
Inequality (1.9), in turn, implies ^$\widehat{\mathbb{P}}_{n}(\tau_{n}\leq T)\leq\widehat{\mathbb{P}}_{n}(\widehat{\tau}_{n}\leq T)$, for $\widehat{\tau}_{n}:=\inf\{t\geq 0,\;\widehat{Y}_{t}^{n}=n-|\overline{\rho}|K_{\mathbb{T}}M_{\gamma}\}$. Finally, since $\widehat{Z}^{n}$ is a ^$\widehat{\mathbb{P}}_{n}$-Brownian motion, we obtain
$\lim_{n\uparrow\infty}\widehat{\mathbb{P}}_{n}(\tau_{n}\leq T)\leq\lim_{n\uparrow\infty}\widehat{\mathbb{P}}_{n}(\widehat{\tau}_{n}\leq T)=\lim_{n\uparrow\infty}\mathbb{P}\left(\sup_{t\in\mathbb{T}}Y_{t}\geq n-|\overline{\rho}|K_{\mathbb{T}}M_{\gamma}\right)=0,$
and it follows that $\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}$ is indeed a true martingale and note that $\lim_{n\uparrow\infty}\widehat{\mathbb{P}}_{n}=\mathbb{Q}$. In the sense that the relation (1.7) holds between the $\mathbb{P}$ and $\mathbb{Q}$ Brownian motions. This concludes the proof of (I) and, as stated in the beginning, the results in (II) then readily follow by direct application of the change of measure theorem.
@@
We now prove (III): the discounted price $\widetilde{S}=S/B$ is a true martingale for $\rho\leq 0$. A straightforward application of Itô’s formula yields under $\mathbb{Q}$
@@
(1.10) $\left\{\begin{array}[]{rl}\displaystyle\frac{\mathrm{d}\widetilde{S}_{t}}{\widetilde{S}_{t}}&=\sqrt{v_{t}}\,\mathrm{d}W^{\mathbb{Q}}_{t},\\ v_{t}&=\displaystyle\psi\left(t,\widehat{Y}_{t}+\int_{0}^{t}k(t,s)\lambda_{s}\mathrm{d}s\right),\\ \widehat{Y_{t}}&=\displaystyle\int_{0}^{t}k(t,s)\mathrm{d}Z^{\mathbb{Q}}_{s}\end{array}\right.$
@@
Define the stopping time $\iota_{n}:=\inf\{t\geq 0,\;\widehat{Y}_{t}=n\}$. For any $t\in\mathbb{T}$, the random function $g(x):=\psi\left(t,x+\int_{0}^{t}k(t,s)\lambda_{s}\mathrm{d}s\right)$ is bounded $\mathbb{Q}$-almost surely on $(-\infty,a]$ by Assumption 1.1(i)-(iii)-(iv), with $\lambda$ in (1.6), and by boundedness of $\gamma$, so that, since $\widetilde{S}$ is a $\mathbb{Q}$-local martingale:
(1.11) $\widetilde{S}_{0}=\mathbb{E}^{\mathbb{Q}}[\widetilde{S}_{T\wedge\iota_{n}}]=\mathbb{E}^{\mathbb{Q}}[\widetilde{S}_{T}1{1}_{\{T<\iota_{n}\}}]+\mathbb{E}^{\mathbb{Q}}[\widetilde{S}_{\iota_{n}}1{1}_{\{T>\iota_{n}\}}].$

### v2

@@ Abstract.
60G15, 60G22, 91G20
@@ Introduction
Rough volatility is a recent paradigm proposed by Gatheral, Jaisson and Rosenbaum [gatheral2022volatility], which has attracted the attention of many academics and practitioners thanks to its numerous attractive properties [bayer2023rough]. Despite some debate about whether volatility should be rough [jaber2022quintic, jaber2024volatility, guyon2023volatility, romer2022empirical], this class of models provides a general framework to analyse both time series of the instantaneous volatility (under the historical measure $\mathbb{P}$) and prices of financial derivatives (under the pricing measure $\mathbb{Q}$). Starting from a rough version of the Bergomi model [Bergomi] under $\mathbb{P}$, Bayer, Friz and Gatheral [BFG15] showed that a deterministic market price of risk preserved its structure under $\mathbb{Q}$ (somehow akin to the Heston model [heston1993closed] specification).
However, the Financial Economics literature has long shown that this market price of risk, monitoring the transition from $\mathbb{P}$ to $\mathbb{Q}$ via Girsanov’s transform, is not constant nor deterministic but instead stochastic. Its estimation has been the source of long academic discussions, outside the scope of the present paper though, and we refer the interested reader to [bansal2002market, carr2016analyzing, chabi2012pricing, duan2014forward, kaido2009inference, mcdonald1984option] for some useful pointers. This of course has serious practical implications for risk management, and [graham2005long] are fascinating sources of information. We focus here on this particular bridge between $\mathbb{P}$ and $\mathbb{Q}$ and show, not surprisingly, that the required stochasticity of the market price of risk unfortunately breaks the structure of the rough Bergomi model under $\mathbb{Q}$. However, we link the Hölder regularity of the volatility process (lower in this class of rough models) with that of the change of measure, and design several specifications making the model tractable under $\mathbb{Q}$. While the rough Bergomi model tracks well the behaviour of the historical volatility, it is less powerful for option prices, especially when considering VIX smiles (which are more or less flat under this model). Our new setup allows for more flexibility there, while preserving the $\mathbb{P}$-tractability of the model.
@@ 1. Rough volatility models and change of measure
Rough volatility models are a natural extension of classical stochastic volatility models. Starting from such a model under the historical measure $\mathbb{P}$, we characterise below its dynamics under martingale measures $\mathbb{Q}$ equivalent to $\mathbb{P}$, which then, by the fundamental theorem of asset pricing, allows for arbitrage-free option pricing. Following [gatheral2022volatility] for example, we consider a rather general class of (rough) stochastic volatility models under $\mathbb{P}$, where the stock price process admits the following dynamics:
@@
starting from $S_{0},v_{0}>0$, over a fixed time interval $\mathbb{T}:=[0,T]$, for $T>0$. Here, $\mu$ is a $(\mathcal{F}_{t})_{t\in\mathbb{T}}$-measurable process, $\psi:\mathbb{T}\times\mathbb{R}\to(0,\infty)$ a continuous function, $\boldsymbol{W}^{\mathbb{P}}=(W^{\mathbb{P}},W^{\mathbb{P},\perp})$ a two-dimensional standard Brownian motion on a given filtered probability space $(\Omega,\mathcal{F},\mathbb{P})$, with $\mathcal{F}:=\mathcal{F}^{W^{\mathbb{P}}}\vee\mathcal{F}^{W^{\mathbb{P},\perp}}$, and $Z^{\mathbb{P}}:=\rho W^{\mathbb{P}}+\overline{\rho}W^{\mathbb{P},\perp}$, where $\rho\in[-1,1]$ and $\overline{\rho}:=\sqrt{1-\rho^{2}}$. We assume that for each $t\in\mathbb{T}$, the kernel $k(t,\cdot)$ is null on $\mathbb{T}\setminus[0,t]$ and $\int_{0}^{\cdot}k(\cdot,s)\mathrm{d}Z^{\mathbb{P}}_{s}$ is a well-defined continuous Gaussian process. In particular, $k(t,\cdot)\in L^{2}([0,t])$ for each $t\in\mathbb{T}$ ensures that the process is well defined, and we refer to [Section 5][alos2001stochastic] or [decreusefond2005stochastic, Section 3] for conditions of continuity criteria. For example, the Gamma kernel, commonly used to model turbulence, and pioneered by Barndorff-Nielsen and Schmiegel [Barndorff-Nielsen], is given by
(1.2) $k(t,s)=(t-s)^{H-\frac{1}{2}}\mathrm{e}^{-\beta(t-s)}1\hskip-5.97508pt{1}_{\{t\geq s\}},\qquad\text{with }H\in(0,1),\quad\beta\geq 0.$
We further introduce the set $\mathbb{F}_{b}$ of $\mathbb{P}$-bounded and $(\mathcal{F}_{t})_{t\in\mathbb{T}}$-progressively measurable processes, namely $X\in\mathbb{F}_{b}$ if there exists a constant $c>0$ such that $\mathbb{P}\left(\sup_{t\geq 0}|X_{t}|\leq c\right)=1$, and recall the Doléans-Dade stochastic exponential of a square integrable continuous process $X$:
$\mathcal{E}(X)_{t}:=\exp{\left(X_{t}-\frac{1}{2}\langle X\rangle_{t}\right)},\qquad\text{for }t\in\mathbb{T}.$
Finally, we introduce a progressively measurable interest rate process $(r_{t})_{t\in\mathbb{T}}$, define the corresponding Sharpe ratio by $\chi_{t}:=\frac{r_{t}-\mu_{t}}{\sqrt{v_{t}}}$, for $t\in\mathbb{T}$ and consider the following assumption:
Assumption 1.1.
The processes $\chi$ and $\gamma$ are $(\mathcal{F}_{t})_{t\in\mathbb{T}}$-adapted with càdlàg paths.
(1.3) $\mathcal{D}^{\gamma}_{t}:=\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}\bigg|_{\mathcal{F}_{t}^{\mathbb{P}}}=\mathcal{E}\left(\int_{0}^{\cdot}\chi_{u}\mathrm{d}W^{\mathbb{P}}_{u}+\int_{0}^{\cdot}\gamma_{u}\mathrm{d}W^{\mathbb{P},{\perp}}_{u}\right)_{t}.$
Proposition 1.2.
Under Assumption 1.1, $\mathcal{D}^{\gamma}$ is a locally integrable $\mathbb{P}$-local martingale on $\mathbb{T}$.
Proof.
Since $W^{\mathbb{P}}$ and $W^{\mathbb{P},\perp}$ are independent Brownian motions (and so locally square-integrable local martingales), thus by [protter2005stochastic, Section II, Theorem 20] the processes $X^{1}:=\int_{0}^{\cdot}\chi_{u}\mathrm{d}W^{\mathbb{P}}_{u}$ and $X^{2}:=\int_{0}^{\cdot}\gamma_{u}\mathrm{d}W^{\mathbb{P},\perp}_{u}$ are locally square-integrable local martingales. The sum of two locally square-integrable local martingales is so itself and hence so is $X:=X^{1}+X^{2}$. Finally, notice that the Radon-Nikodym derivative  $\mathcal{D}^{\gamma}$ defined in (1.3) is precisely the stochastic exponential of $X$ and so a non-negative local martingale itself. This implies that $\mathcal{D}^{\gamma}$ is a positive super-martingale which together with the fact that $\mathcal{D}^{\gamma}_{0}=1$ yields $\mathcal{D}^{\gamma}_{t}\in L^{1}$ for all $t\in\mathbb{T}$. ∎
@@
    $K_{\mathbb{T}}>0$ such that $\sup_{t\in\mathbb{T}}\left\{\int_{0}^{t}k(t,u)\lambda_{u}\mathrm{d}u\right\}\leq K_{\mathbb{T}}$, $\mathbb{P}$-almost surely, where $\lambda$ denotes the market price of volatility risk defined by
    (1.4) $\lambda_{t}:=\rho\chi_{t}+\overline{\rho}\;\gamma_{t};$
@@
    The Radon-Nikodym derivative satisfies $\mathbb{E}[\mathcal{D}^{\gamma}_{t}]=1$ for all $t\in\mathbb{T}$.
@@
Note that Assumption 1.3(ii) also implies that, for all $n\in\mathbb{N}$,
(1.5) $\sup_{t\in\mathbb{T}}\left\{\int_{0}^{t}k(t,u)\lambda_{u}\mathrm{d}u\right\}\leq K_{\mathbb{T}},\qquad{\widehat{\mathbb{P}}_{n}}\text{-almost surely},$
with
(1.6) $\frac{\mathrm{d}\widehat{\mathbb{P}}_{n}}{\mathrm{d}\mathbb{P}}\bigg|_{\mathcal{F}_{t}}:=\mathcal{E}\left(\int_{0}^{\cdot}\chi_{u}\mathrm{d}W^{\mathbb{P}}_{u}+\int_{0}^{\cdot}\gamma_{u}\mathrm{d}W^{\mathbb{P},{\perp}}_{u}\right)_{t\wedge\tau_{n}}\quad\text{and}\quad\tau_{n}:=\inf\{t\geq 0,\;Y_{t}=n\}.$
Indeed, given an event $A$ such that $\mathbb{P}(A)=1$, Cauchy-Schwarz inequality yields
$1$ $=\mathbb{P}(A)=\mathbb{E}[\mathbb{1}_{A}]=\widehat{\mathbb{E}}_{n}[\mathbb{1}_{A}\mathcal{E}(X)^{-1}]\leq\widehat{\mathbb{E}}_{n}[\mathbb{1}_{A}]^{\frac{1}{2}}\widehat{\mathbb{E}}_{n}[\mathcal{E}(X)^{-2}]^{\frac{1}{2}}=\widehat{\mathbb{P}}_{n}(A)^{\frac{1}{2}}{\mathbb{E}}[\mathcal{E}(X)^{-2}\mathcal{E}(X)]^{\frac{1}{2}}$
$=\widehat{\mathbb{P}}_{n}(A)^{\frac{1}{2}}{\mathbb{E}}[\mathcal{E}(X)^{-1}]^{\frac{1}{2}}=\widehat{\mathbb{P}}_{n}(A)^{\frac{1}{2}}{\mathbb{E}}\left[\mathcal{E}(-X)\exp\left(\int_{0}^{t\land\tau_{n}}\chi_{u}^{2}+\gamma_{u}^{2}\mathrm{d}u\right)\right]^{\frac{1}{2}}\leq\widehat{\mathbb{P}}_{n}(A)^{\frac{1}{2}},$
where in the last step we have exploited the fact that $\mathcal{E}(-X)$ is a non-negative supermartingale since X-X a continuous local martingale. Thus, we have proved $\widehat{\mathbb{P}}_{n}(A)=1$. An analogous argument shows that the same holds for $\mathbb{Q}$.
Remark 1.5.
Assumption 1.3(iv) is guaranteed under different sets of stronger assumptions on $\gamma$, $\mu$, $\chi$, $r$ and $v$, in particular
  • if $\mathbb{E}\left[\exp\left\{\frac{1}{2}\int_{0}^{T}|X_{t}|^{2}\mathrm{d}t\right\}\right]<\infty$, namely $X$ satisfy the Novikov condition;
  • if Assumption 1.1 and 1.3(i)-(ii) hold and processes $\mu,\gamma,r$ belong to $\mathbb{F}_{b}$, as detailed in Appendix A.
Proposition 1.2 justifies the use of a Doléans-Dade exponential in the definition of $\mathcal{D}^{\gamma}$.
@@
Under Assumptions 1.1 and 1.3, the following hold:
@@
    with $S_{0},v_{0}>0$, and $\lambda$ is the market price of volatility risk in (1.4) and where $W^{\mathbb{Q}}$ and $Z^{\mathbb{Q}}$ are $\mathbb{Q}$-Brownian motions defined as
    (1.8) $W^{\mathbb{Q}}:=\displaystyle W^{\mathbb{P}}+\int_{0}^{\cdot}\chi_{u}\mathrm{d}u\qquad\text{and}\qquad Z^{\mathbb{Q}}:=Z^{\mathbb{P}}+\int_{0}^{\cdot}\lambda_{u}\mathrm{d}u;$
@@
    the discounted stock price $\widetilde{S}:=\frac{S}{B}$ with $\mathrm{d}B_{t}=r_{t}B_{t}\mathrm{d}t$, $B_{0}=1$, is a true $\mathbb{Q}$-martingale.
@@
To satisfy the no-arbitrage conditions, the change of measure for $W^{\mathbb{P}}$ is constrained by the martingale restriction on the discounted spot dynamics, while the Brownian motion $Z^{\mathbb{P}}$ gives freedom to the model and makes the market incomplete by the free choice of the process $\gamma$. Consequently, the change of measure from $\mathbb{P}$ to $\mathbb{Q}$ and the corresponding Radon-Nikodym derivative directly follow from Girsanov’s Theorem via (1.3), provided that $\mathcal{D}^{\gamma}_{t}\in L^{1}$ and $\mathcal{D}^{\gamma}$ is a true martingale. Thus, once we have shown (I), then (II) automatically follows. By Proposition 1.2, $\mathcal{D}^{\gamma}_{t}\in L^{1}$ and, being a non-negative local martingale, is a super-martingale, and a true martingale on $\mathbb{T}$ if and only if $\mathbb{E}\left[\mathcal{D}^{\gamma}_{T}\right]=1$. This is guaranteed by Assumption 1.3(iv), hence (I) holds, and therefore (II) as well.
@@
We now prove (III): the discounted price $\widetilde{S}=\frac{S}{B}$ is a true martingale for $\rho\leq 0$. Itô’s formula under $\mathbb{Q}$ yields
@@
(1.9) $\left\{\begin{array}[]{rl}\displaystyle\frac{\mathrm{d}\widetilde{S}_{t}}{\widetilde{S}_{t}}&=\sqrt{v_{t}}\,\mathrm{d}W^{\mathbb{Q}}_{t},\\ v_{t}&=\displaystyle\psi\left(t,\widehat{Y}_{t}+\int_{0}^{t}k(t,s)\lambda_{s}\mathrm{d}s\right),\\ \widehat{Y_{t}}&=\displaystyle\int_{0}^{t}k(t,s)\mathrm{d}Z^{\mathbb{Q}}_{s}.\end{array}\right$.
@@
Define the stopping time $\iota_{n}:=\inf\{t\geq 0,\;\widehat{Y}_{t}=n\}$. For any $t\in\mathbb{T}$, the random function $g(x):=\psi\left(t,x+\int_{0}^{t}k(t,s)\lambda_{s}\mathrm{d}s\right)$ is bounded $\mathbb{Q}$-almost surely on $(-\infty,a]$ by Assumption 1.3(i)-(ii), with $\lambda$ in (1.4), so that, since $\widetilde{S}$ is a $\mathbb{Q}$-local martingale:
(1.10) $\widetilde{S}_{0}=\mathbb{E}^{\mathbb{Q}}[\widetilde{S}_{T\wedge\iota_{n}}]=\mathbb{E}^{\mathbb{Q}}[\widetilde{S}_{T}1\hskip-5.97508pt{1}_{\{T<\iota_{n}\}}]+\mathbb{E}^{\mathbb{Q}}[\widetilde{S}_{\iota_{n}}1\hskip-5.97508pt{1}_{\{T>\iota_{n}\}}].$